# CRYSP: a Total-Body PET based on cryogenic cesium iodide crystals

### v1

@@ Abstract
Total Body PET (TBPET) scanners have recently demonstrated the ability to significantly reduce both acquisition time and the administered radioactive dose, thanks to their increased sensitivity. However, their widespread adoption is limited by the high costs associated with the current available systems. In this context, pure cesium iodide (CsI) monolithic crystals, given their much lower cost compared with currently used rare-earth crystals, offer a promising solution to improve accessibility. When operated at cryogenic temperatures (approximately 100 K), CsI crystals exhibit one of the highest light outputs among inorganic crystals, around $10^{5}$ photons/MeV. This results in energy resolution below 7%, millimeter-scale spatial resolution (including accurate depth-of-interaction determination), and coincidence time resolution at the nanosecond level, despite their relatively slow scintillation decay time. This study demonstrates that a TBPET scanner based on cryogenic CsI monolithic crystals might have the potential to deliver high-performance imaging at a reduced cost compared with conventional systems, paving the way for broader deployment and accessibility.
Index Terms:
@@ I Introduction
Figure 1: Increase in component costs for a 70-cm, 100-cm, 140-cm and 200-cm system versus a system with 20 cm axial length; the y-axis is the system cost in relative units. Data taken from ref.. [1].
The last decade has witnessed the development of Total Body PET (TBPET) scanners with axial fields-of-view (AFOV) between 3 and 7 times that of conventional PET (CPET) apparatus. Such devices detect a large fraction of the emitted photons, thus dramatically increasing sensitivity, and allowing for simultaneous dynamic acquisition from all tissues of interest [2, 1]. TBPETs also offer the flexibility of optimizing signal-to-noise ratio (SNR), reducing the acquisition time, or decreasing the amount of administered radioactive dose.
Prime examples of such TBPETs are the United Imaging uEXPLORER[3], and Siemens Biograph Vision Quadra [4]. The AFOV of uEXPLORER is 194 cm, while the AFOV of the smaller Quadra is 106 cm, to be compared with an AFOV of 20-30 cm typical of a CPET (e.g., the new generation General Electric Omni Legend 32 [5] has an AFOV of 32 cm). Both TPBET systems use arrays of pixelated lutetium-yttrium oxyorthosilicate (LYSO) crystals, read out by SiPMs and custom electronics, with very high sensitivity and excellent time-of-flight (TOF) resolution. The approximate increase in cost of the uEXPLORER (Quadra) compared with a typical CPET is roughly 8 (4) times (see fig. 1), well above what can be afforded by most hospitals.
Figure 2: The geometrical acceptance for a point-like source placed at the axial center, a point-like source placed 10 cm from one end of the scanner, and line sources of 10 cm, 100 cm and 200 cm length, as a function of the axial length. It is calculated in the transverse centre of a PET scanner with a diameter of 80 cm.
The bulk of the cost in both systems comes from the scintillator, followed by the cost of the SiPMs and electronics. Of the three items, the cost of SiPMs, in large quantities, is roughly proportional to the total surface of chips deployed (not to the size of the chip), and thus proportional to the size and coverage of the scanner. However, the cost of SiPMs decreases about 5-10 % annually, with current prices about one third of those ten years ago. Thus their relative contribution to the total cost of future TPBETs will steadily decrease with time. Instead, the cost of crystals and that of the readout electronics has remained stable, and in the case of LYSO, significantly expensive. Thus, the first priority of any putative low-cost TBPET is to find significantly cheaper alternatives to LYSO. Notice that, in addition to the high cost, LYSO is a rare earth whose production and manufacturing is heavily concentrated. Last but not least, the cost of the readout electronics can be decreased for large scale production developing specific ASICs, and, importantly, reducing the number of electronics channels.
@@
The choice for a PET scanner of 100 cm length is also motivated by the approximate axial distribution of whole-body organs of interest (e.g., brain to pelvis) for typical human height. The figure shows that, with an extended source of 100 cm, one gains in geometrical acceptance from 50% (100 cm scanner), to about 80% (200 cm scanner), a modest increased compared with cost. Thus, an AFOV of about 100 cm as implemented by the Quadra appears as a good compromise between cost and performance and a reasonable target for any TPBET proposal.
Figure 3: (a) Compton scatter detection in PET. In Compton scattering a photon (in yellow) undergoes an interaction within the patient’s body where it changes direction and loses energy. Thus, the PET scanner reconstructs a wrong LOR (in red), affecting the signal-to-background ratio of the reconstructed image. Adapted from ref. [6]. (b) The residual energy $E_{\gamma^{\prime}}$ of a 511-keV incoming photon after Compton scattering with a given angle $\theta$.
Figure 4: Mean parallax error $\Delta\theta$, defined as the angular distance between the ideal LOR and the reconstructed LOR, as a function of the LOR axial angle. The result was obtained by simulating a PET scanner with an axial length of 1 m and a diameter of 0.8 m, made of $3\times 3\times 22.4$ mm3 LYSO pixels.
@@
Conventional PET scanners are made of pixelated detectors, with crystals of small transverse dimensions and a thickness of typically 2 radiation lengths (e.g., $3.2\times 3.2\times 20$ mm3 for the Quadra). For a gamma impinging at angles normal to the crystal surface (small transaxial angles), the resolution in the transaxial coordinates depends only on the transverse size of the pixel, and the resolution on the depth of impact (d.o.i) depends on the pixel thickness. However, for large axial angles parallax error increases (see fig. 4), due to the lack of d.o.i measurement in pixelated crystals [7, 8].
@@
Although techniques exist to compensate for CS and PE [9], their overall effect is to degrade the contribution to image reconstruction of gammas emitted at high angles and therefore to partially negate the increase in sensitivity afforded by the longer axial coverage of a TPBET.
@@
In this paper, we propose a new type of TPBET scanner called CRYSP [^1] The acronym stands simultaneously for CRYStal PET and for CRYogenic ceSium PET.. CRYSP is based on large monolithic crystals of CsI crystals made of pure cesium iodide (CsI) operating at cryogenic temperatures. The light output of these crystals is very large (about 105 photons/MeV [10, 11, 12]), which translates into excellent energy resolution. This performance, combined with the monolithic form factor, allows to achieve a millimeter-scale spatial resolution in all three dimensions, minimizing the PE. As we will show in the next sections, these features allow to achieve performances that are comparable to the ones of current state-of-the-art TPBET scanners, at a significantly reduced cost.
@@
This paper is organized as follows. Section II discusses the potential of pure, cryogenic CsI as scintillator material for PET, and how the choice of this material leads naturally to monolithic crystals. Section III describes the CRYSP baseline design and section IV focuses on the dedicated electronics. The modelization of the apparatus using the Geant4 simulation framework [13] is detailed in section V. Section VI discusses our strategy to reconstruct the gamma impact coordinates using convolutional neural networks (CNNs). In section VII we discuss the main parameters illustrating CRYSP performance following the NEMA guidelines (e.g., sensitivity, NECR, spatial resolution). Section VIII offers an evaluation of the apparatus performance using a modified Jaszczak phantom, comparing it with a reference detector based on pixelated LYSO crystals. Conclusions are presented in section IX.
@@
TABLE I: Properties of scintillating materials commonly used for PET, compared with scintillators based on cesium iodide. Data for LYSO, BGO and CsI(Tl) have been obtained from vendor [14]. Values for CsI were measured in ref. [12]. Cost is normalized by one radiation length $X_{0}$. The range in prices depends on the quality of the material and the manufacturing process, with pixelated blocks being more expensive that monolithic crystals for a fixed volume.
@@
II CsI as scintillator material for PET.
Table I shows the properties of cesium iodide compared with other crystals commonly used for PET. Pure CsI at ambient temperature has been adopted in particle physics for experiments set in high-rate and high-radiation environments, given its fast decay time and radiation hardness [15, 16]. However, at a first glance, its low light yield (5 photons/keV), UV emission, average radiation length and density make it a sub-optimal candidate for PET.
Cooling pure CsI to cryogenic temperatures, approximately 100 K, causes its light emission spectrum to shift towards the near ultraviolet range, moving from around 310 nm to 340 nm [17]. This process also increases its light yield by a factor of 20, to about 100 photons/keV [10, 11, 18]. Unfortunately, its decay time also increases, approximately by a factor of 50, from 15 ns to 800 ns. In a PET scanner, the coincidence time resolution (CTR) between pairs of crystals is determined by the material’s emission time profile and amount of light detected. An approximate figure of merit is given by [19]:
@@
Thus, since the time resolution at room temperature is at the nanosecond level [20], effective coincidence timing should be possible also at 100 K, in spite of the long decay time, as we will discuss further in section VII.
In a previous paper [12] we measured the energy resolution and coincident time resolution (CTR) for a pair of pure CsI crystals at cryogenic temperatures read out by solid state photosensors (Hamamatsu MPPCs). At $T=104$ K, an energy resolution better than 7% FWHM and a coincidence time resolution lower than 2 ns FWHM have been achieved. While these are already very promising results, we believe there is still room for improvement, in particular when using large monolithic crystals, SiPMs with higher PDE at 350 nm and optimized electronics. In fact, the intrinsic resolution of cryogenic CsI at 511 keV is approximately 5% [21], and one of the crystals showed a resolution of 6.3%, which we will take as default for the studies presented here. For the CTR, we use a conservative estimate of 1.5 ns, based on the 1.84 ns obtained with a tabletop setup.
III CRYSP
The basic building block of CRYSP is a monolithic crystal, of dimensions $48\times 48\times 37~{}$ mm3, where the crystal thickness corresponds to 2 radiation lengths. All crystal surfaces are polished, and five of the six faces are wrapped in PTFE to ensure a reflectivity $>95\%$. The wrapping adds about 2 mm to the crystal transverse dimensions. Full wrapping with a good reflector is a must for energy resolution, and in the case of CsI, the relatively hard emission wavelength limits the potential reflectors, with PTFE being the best option [22]. Each crystal is read out by an array of $8\times 8$ Hamamatsu S14160-6050HS MPPCs [23]. These photosensors have an active area of $6\times 6~{}$ mm2 and a photon detection efficiency of approximately 38% at the emission wavelength of cryogenic CsI (see fig. 5).
Figure 5: CsI scintillation spectrum at room and cryogenic temperatures [17] (in black, left axis), compared with the photon detection efficiency of the Hamamatsu S14160-6050HS MPPC [23] (in red, right axis).
Thus, an interaction in the crystal results in an image of $8\times 8$ pixels, where the intensity in each pixel is determined by the photoelectrons recorded by each SiPM (fig. 6). System specifications are detailed in table II.
TABLE II: CRYSP1M scanner specifications.
@@ III CRYSP
Figure 6: Charge collected by the SiPM matrix coupled to a monolithic crystal for two different interactions, one close to the SiPMs plane ($z=-12.20$ mm) and one near the crystal entrance ($z=17.30$ mm). The red dot corresponds to the simulated interaction point. The difference between the two patterns is exploited by a neural network to reconstruct the d.o.i. (here the $z$ coordinate).
For definiteness, we consider a CRYSP scanner with dimensions similar to those of the Quadra (CRYSP1M). This corresponds to an effective diameter of $\sim 774~{}$mm and an effective length of $\sim 1024~{}$mm, which translates to 20 rings, where each ring is comprised of 48 crystals of $50\times 50\times 38$ mm3, including wrapping. Thus, CRYSP1M deploys a total of 960 crystals and 61,440 channels (compared with the 243,200 of the Quadra).
@@
Figure 7 shows a conceptual design. The scanner is hosted in a cryostat. The inner volume of the cryostat is filled with liquid nitrogen that bathes the crystals and the SiPMs, keeping them at approximately 78 K. This implies that dark current in the SiPMs is reduced, and its effect negligible, given the very high yield of the scintillator. The outer volumes of the cryostat are at vacuum and provide thermal isolation.
The total estimated heat loss of the cryostat is 20 W. Out of these losses, 70% are associated with the inner instrumentation of the SIPMs that have a direct thermal path between the inner and outer vessels. The other 30% are radiation losses assuming a 10 layer Multi-layer Insulation of aluminized Mylar surrounding the inner cryostat. The volume of liquid nitrogen evaporated per hour due to 20 W of heat loss is  0.45 L/h making a yearly consumption of  3942 L/year. The typical costs of liquid nitrogen are below 0.6 €/L in facilities with on site liquid nitrogen storage tanks, therefore the associated running costs will be close to 2,000 €/year.
Figure 8: Detail of the scanner showing the connection between four SiPM boards (in green), coupled to four monolithic crystals (in purple), and the exterior. In this conceptual design, the feedthroughs consist of rigid PCBs acting as feedthroughs connected to flexible cables in the vacuum region to allow for the inner connection the cables and the mounting of the flanges. A third, thinner vacuum chamber (shown at the bottom in this illustration) separates the liquid nitrogen-filled chamber from the external environment near the patient.
Each SiPM board is connected to a simple feedthrough (e.g., a potted rigid PCB circuit, see fig. 8), that allows extracting the signal outside the cryostat, where the readout electronics are placed. The connection in between the inner and external feedthrough will be done by flexible cables allowing for the connection and mounting of the different flanges. This arrangement has already been validated in similar environments and does not represent a particular technical challenge [24].
When a gamma interacts in a crystal, the readout ASIC digitizes the charge in each of the 64 channels and sends $64\times 2$ numbers to an FPGA, corresponding to the integrated charge and the timestamp value, as described in section IV. The FPGA computes the time of interaction assigned to the crystal (e.g., by taking an OR of the individual timestamps), as well as the total energy and the interaction vertex. The latter is achieved by passing the digitized signals (the $8\times 8$ pixels comprising the “image” provided by the crystal) in real-time through a pre-trained CNN running on the FPGA itself (see section VI). Thus, the information provided by each crystal is reduced to five numbers: time, energy and the three coordinates of the interaction vertex.
@@
Although the CTR is good enough to resolve coincidences, a fraction of events will still be lost or misreconstructed because of pile-up (e.g., events in which a second gamma interacts in the crystal before the integration of the energy of the first event is over). A dedicated pile-up processor is described in section IV-B and its effect is quantified in section VII-D.
IV Electronics for CRYSP

### v2

@@ Abstract
Objective. Total Body PET (TBPET) scanners have the potential to substantially reduce both acquisition time and administered radiation dose, owing to their high sensitivity. However, their widespread clinical adoption is hindered by the high cost of currently available systems. This work explores the use of pure cesium iodide (CsI) monolithic crystals operated at cryogenic temperatures as a cost-effective alternative to rare-earth scintillators for TBPET. Approach. We investigate the performance of pure CsI crystals operated at cryogenic temperatures ($\sim$100 K), where they achieve a light yield of approximately $10^{5}$ photons/MeV. The implications for energy resolution, spatial resolution (including depth-of-interaction capability), and timing performance are assessed, with a view toward their integration into a TBPET system. Main results. Cryogenic CsI crystals demonstrated energy resolution below 7% and coincidence time resolution at the nanosecond level, despite their relatively slow scintillation decay time. A Monte Carlo simulation of monolithic CsI crystals shows that a millimeter-scale spatial resolution in all three dimensions can be obtained. These characteristics indicate that high-performance PET imaging is achievable with this technology. Significance. A TBPET scanner based on cryogenic CsI monolithic crystals could combine excellent imaging performance with significantly reduced detector costs, enabling broader accessibility and accelerating the adoption of TBPET in both clinical and research settings.
keywords:
@@ 1 Introduction
Figure 1: Increase in component costs for a 70-cm, 100-cm, 140-cm and 200-cm system versus a system with 20 cm axial length; the y-axis is the system cost in relative units. Data taken from ref. [vandenberghe2020state].
The last decade has witnessed the development of Total Body PET (TBPET) scanners with axial fields-of-view (AFOV) between 3 and 7 times that of conventional PET (CPET) apparatus. Such devices detect a large fraction of the emitted photons, thus dramatically increasing sensitivity, and allowing for simultaneous dynamic acquisition from all tissues of interest [vandenberghe2020state, nadig2022hybrid]. TBPETs also offer the flexibility of optimizing signal-to-noise ratio, reducing the acquisition time, or decreasing the amount of administered radiation dose.
Prime examples of such TBPETs are the United Imaging uEXPLORER [cherry2018total], and Siemens Biograph Vision Quadra [prenosil2022performance]. The AFOV of uEXPLORER is 194 cm, while the AFOV of the smaller Quadra is 106 cm, to be compared with an AFOV of 20-30 cm typical of a CPET (e.g., the new generation General Electric Omni Legend 32 [yamagishi2023performance] has an AFOV of 32 cm). Both TBPET systems use arrays of pixelated lutetium-yttrium oxyorthosilicate (LYSO) crystals, read out by SiPMs and custom electronics, with very high sensitivity and excellent time-of-flight (TOF) resolution. The approximate increase in cost of the uEXPLORER (Quadra) compared with a typical CPET is roughly 8 (4) times (see fig. 1), well above what can be afforded by most hospitals.
Figure 2: The geometrical acceptance for a point-like source placed at the axial center, a point-like source placed 10 cm from one end of the scanner, and line sources of 10 cm, 100 cm and 200 cm length, as a function of the axial length. It is calculated in the transverse center of a PET scanner with a diameter of 80 cm.
The bulk of the cost in both systems comes from the scintillator, followed by the cost of the SiPMs and electronics. Of the three items, the cost of SiPMs, in large quantities, is roughly proportional to the total surface of chips deployed (not to the size of the chip), and thus proportional to the size and coverage of the scanner. However, the cost of SiPMs decreases about 5-10 % annually, with current prices about one third of those ten years ago. Thus their relative contribution to the total cost of future TBPETs will steadily decrease with time. Instead, the cost of crystals and that of the readout electronics has remained stable, and in the case of LYSO, significantly expensive. Thus, the first priority of any putative low-cost TBPET is to find significantly cheaper alternatives to LYSO. Notice that, in addition to the high cost, LYSO is a rare earth whose production and manufacturing is heavily concentrated. Last but not least, the cost of the readout electronics can be decreased for large scale production developing specific ASICs, and, importantly, reducing the number of electronics channels.
@@
The choice for a PET scanner of 100 cm length is also motivated by the approximate axial distribution of whole-body organs of interest (e.g., brain to pelvis) for typical human height. The figure shows that, with an extended source of 100 cm, one gains in geometrical acceptance from 50% (100 cm scanner), to about 80% (200 cm scanner), a modest increased compared with cost. Thus, an AFOV of about 100 cm as implemented by the Quadra appears as a good compromise between cost and performance and a reasonable target for any TBPET proposal.
Figure 3: (a) Compton scatter detection in PET. In Compton scattering a photon (in yellow) undergoes an interaction within the patient’s body where it changes direction and loses energy. Thus, the PET scanner reconstructs a wrong LOR (in red), affecting the signal-to-background ratio of the reconstructed image. (b) The residual energy $E_{\gamma^{\prime}}$ of a 511-keV incoming photon after Compton scattering with a given angle $\theta$.
@@
Conventional PET scanners are made of pixelated detectors, with crystals of small transverse dimensions and a thickness of typically 2 radiation lengths (e.g., $3.2\times 3.2\times 20$ mm3 for the Quadra). For a gamma impinging at angles normal to the crystal surface (small transaxial angles), the resolution in the transaxial coordinates depends only on the transverse size of the pixel, and the resolution on the depth-of-interaction (d.o.i.), which depends on the pixel thickness, does not significantly affect the reconstruction. While d.o.i. information can be incorporated into pixelated detectors through dedicated schemes (such as dual-ended readout, phoswich configurations, or depth-segmented crystals), the standard pixelated blocks used in current TBPET systems do not provide this information, increasing the PE at large angles [schmall2016parallax, surti2020total].
@@
Although techniques exist to compensate for CS and PE [zaidi2007scatter], their overall effect is to degrade the contribution to image reconstruction of gammas emitted at high angles and therefore to partially negate the increase in sensitivity afforded by the longer axial coverage of a TBPET.
@@
In this paper, we propose a new type of TBPET scanner called CRYSP [^1] The acronym stands simultaneously for CRYStal PET and for CRYogenic ceSium PET.. CRYSP is based on large monolithic crystals made of pure cesium iodide (CsI) operating at cryogenic temperatures. The light output of these crystals is very high (about 105 photons/MeV [mikhailik2015luminescence, Lewis:2021cjv]), which translates into excellent energy resolution, below 7% at 511 keV [Soleti:2024flw]. This performance, combined with the monolithic form factor, allows to achieve a millimeter-scale spatial resolution in all three dimensions, minimizing the PE. As we will show in the next sections, these features allow to achieve performances that are comparable to the ones of current state-of-the-art TBPET scanners, at a significantly reduced cost.
@@
This paper is organized as follows. Section 2.1 discusses the potential of pure, cryogenic CsI as scintillator material for PET, and how the choice of this material leads naturally to monolithic crystals. Section 2.2 describes the CRYSP baseline design and section 2.3 focuses on the dedicated electronics. The modelization of the apparatus using the Geant4 simulation framework [GEANT4:2002zbu] is detailed in section 2.4. Section 2.5 discusses our strategy to reconstruct the gamma impact coordinates using convolutional neural networks (CNNs). The main parameters illustrating CRYSP performance following the NEMA guidelines (e.g., sensitivity, NECR, spatial resolution), compared with a reference detector based on pixelated LYSO crystals, are discussed throughout sections 3.1 to 3.4. Section 3.5 offers an evaluation of the two apparatus imaging performances using a modified Jaszczak phantom. A discussion of the implications of our findings on the development of TBPET scanners is included in section 4. Conclusions are presented in section 5.
@@
Table 1: Properties of scintillating materials commonly used for PET, compared with scintillators based on cesium iodide. Data for LYSO, BGO and CsI(Tl) have been obtained from vendor [Luxium] and are intended at room temperature. Values for CsI were measured in ref. [Soleti:2024flw]. Cost is normalized by one radiation length $X_{0}$. The range in prices depends on the quality of the material and the manufacturing process, with pixelated blocks being more expensive that monolithic crystals for a fixed volume.
@@
2 Methods and materials
2.1 CsI as a scintillator material for PET
Table 1 shows the properties of cesium iodide compared with other crystals commonly used for PET. Pure CsI at ambient temperature has been adopted in particle physics for experiments set in high-rate and high-radiation environments, given its fast decay time and radiation hardness [Atanov:2016hoz, Doroshenko:2005gd]. However, at a first glance, its low light yield (5 photons/keV), UV emission, average radiation length and density make it a sub-optimal candidate for PET.
Cooling pure CsI to cryogenic temperatures, approximately 100 K, causes its light emission spectrum to shift towards the near ultraviolet range, moving from around 310 nm to 350 nm [amsler2002temperature]. This process also increases its light yield by a factor of 20, to about 100 photons/keV [mikhailik2015luminescence, Lewis:2021cjv, Ding:2023pqe]. Unfortunately, its decay time also increases, approximately by a factor of 50, from 15 ns to 800 ns. In a PET scanner, the coincidence time resolution (CTR) between pairs of crystals is determined by the material’s emission time profile and amount of light detected. An approximate figure of merit is given by [conti2009comparison]:
@@ 2.1 CsI as a scintillator material for PET
Thus, since the time resolution at room temperature is at the nanosecond level [kubota1988new], effective coincidence timing should be possible also at 100 K, in spite of the long decay time, as we will discuss further in section 3.
In a previous paper [Soleti:2024flw] we measured the energy resolution and coincident time resolution (CTR) for a pair of pure CsI crystals at cryogenic temperatures read out by solid state photosensors (Hamamatsu MPPCs). At $T=104$ K, an energy resolution better than 7% FWHM and a coincidence time resolution lower than 2 ns FWHM have been achieved. While these are already very promising results, we believe there is still room for improvement, in particular when using large monolithic crystals, SiPMs with higher PDE at 350 nm and optimized electronics. In fact, the intrinsic resolution of cryogenic CsI at 511 keV is approximately 5% [moszynski2005energy], and one of the crystals showed a resolution of 6.3%, which we will take as default for the studies presented here. For the CTR, we use a conservative estimate of 1.5 ns, based on the 1.84 ns obtained with the tabletop setup of ref. [Soleti:2024flw].
2.2 CRYSP
The basic building block of CRYSP is a monolithic crystal, of dimensions $48\times 48\times 37$~ mm3, where the crystal thickness corresponds to 2 radiation lengths. All crystal surfaces are polished, and five of the six faces are wrapped in PTFE to ensure a reflectivity $95\%$. The wrapping adds about 2 mm to the crystal transverse dimensions. Full wrapping with a good reflector is a must for energy resolution, and in the case of CsI, the relatively hard emission wavelength limits the potential reflectors, with PTFE being the best option [janecek2012reflectivity]. Each crystal is read out by an array of $8\times 8$ Hamamatsu S14160-6050HS MPPCs [Hamamatsu]. These photosensors have an active area of $6\times 6$~ mm2 and a photon detection efficiency of approximately 38% at the emission wavelength of cryogenic CsI [Hamamatsu]. Low-temperature SiPM operation is already well established in large-scale particle physics experiments [Matteucci:2025msy, DUNE:2021hwx, Ieki:2018pbf]. At 77 K, the dark count rate is reduced by several orders of magnitude, afterpulsing and optical cross-talk are suppressed, and the gain remains stable after compensating for the decrease of breakdown voltage with temperature ($\approx 50$ mV/K) [NepomukOtte:2016ktf, Hamamatsu]. The photon detection efficiency in the blue–UV region is mostly unaffected at cryogenic temperatures, making SiPMs well suited for the scintillation spectrum of CsI [Iwai:2019scy, Curras-Rivera:2025clp].
Figure 4: Charge collected by the SiPM matrix coupled to a monolithic crystal for two different interactions, one close to the SiPMs plane (z=12.20z=-12.20 mm) and one near the crystal entrance ($z=17.30$ mm). The red dot corresponds to the simulated interaction point. The difference between the two patterns is exploited by a neural network to reconstruct the d.o.i. (here the $z$ coordinate).
Thus, an interaction in the crystal results in an image of $8\times 8$ pixels, where the intensity in each pixel is determined by the photoelectrons recorded by each SiPM (fig. 4).
Table 2: CRYSP1M scanner specifications.
@@ 2.2 CRYSP
For definiteness, we consider a CRYSP scanner with dimensions similar to those of the Quadra (CRYSP1M). This corresponds to an effective diameter of $\sim 774$~$m$ and an effective length of $\sim 1024$~$m$, which translates to 20 rings, where each ring is comprised of 48 crystals of $50\times 50\times 38$ mm3, including wrapping. Thus, CRYSP1M deploys a total of 960 crystals and 61,440 channels (compared with the 243,200 of the Quadra). System specifications are detailed in table 2.
The CRYSP scanner is enclosed within a double‐walled cryostat filled with liquid nitrogen (LN2) that maintains the crystals and SiPMs at approximately 78 K, as shown in figure 5. The estimated thermal load of 365 W corresponds to an LN2 evaporation rate of 7.6 L/h, resulting in an annual consumption below 70,000 L ($\approx 30,000$ €/year). The inner and outer vessels are separated by vacuum layers with multilayer insulation, and all SiPM signals are extracted via compact PCB feedthroughs validated in similar cryogenic environments [ferrario2022status, PETALO:2024byf]. These mechanical and thermal parameters define the stable operating temperature assumed in simulation and demonstrate the feasibility of continuous cryogenic operation.
@@
When a gamma interacts in a crystal, the readout ASIC digitizes the charge in each of the 64 channels and sends $64\times 2$ numbers to an FPGA, corresponding to the integrated charge and the timestamp value, as described in section 2.3. The FPGA computes the time of interaction assigned to the crystal (e.g., by taking an OR of the individual timestamps), as well as the total energy and the interaction vertex. The latter is achieved by passing the digitized signals (the $8\times 8$ pixels comprising the “image” provided by the crystal) in real-time through a pre-trained CNN running on the FPGA itself (see section 2.5). Thus, the information provided by each crystal is reduced to five numbers: time, energy and the three coordinates of the interaction vertex.
@@
Although the CTR is good enough to resolve coincidences, a fraction of events will still be lost or misreconstructed because of pile-up (e.g., events in which a second gamma interacts in the crystal before the integration of the energy of the first event is over). A dedicated pile-up processor is described in section 2.3.2 and its effect is quantified in section 3.4.
2.3 Electronics for CRYSP